One last question I am stuck on. Please help:
x(9x+7)=4x^2 -4x+36

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

x(9x+7)=4x^2 -4x+36
9x^2+7x = 4x^2 -4x+36
5x^2 + 11x - 36 = 0
Can you do the rest?

Answer by ewatrrr(24785)

x(9x+7)=4x^2 -4x+36
9x^2 + 7x = 4x^2 -4x+36
5x^2 + 11x - 36 = 0
(5x - 9)(x +4) = 0, x = -4, 9/5